### v1

# RoboMamba: Multimodal State Space Model for Efficient Robot Reasoning and Manipulation

Abstract
A fundamental objective in robot manipulation is to enable models to comprehend visual scenes and execute actions. Although existing robot Multimodal Large Language Models (MLLMs) can handle a range of basic tasks, they still face challenges in two areas: 1) inadequate reasoning ability to tackle complex tasks, and 2) high computational costs for MLLM fine-tuning and inference. The recently proposed state space model (SSM) known as Mamba demonstrates promising capabilities in non-trivial sequence modeling with linear inference complexity. Inspired by this, we introduce RoboMamba, an end-to-end robotic MLLM that leverages the Mamba model to deliver both robotic reasoning and action capabilities, while maintaining efficient fine-tuning and inference. Specifically, we first integrate the vision encoder with Mamba, aligning visual data with language embedding through co-training, empowering our model with visual common sense and robot-related reasoning. To further equip RoboMamba with action pose prediction abilities, we explore an efficient fine-tuning strategy with a simple policy head. We find that once RoboMamba possesses sufficient reasoning capability, it can acquire manipulation skills with minimal fine-tuning parameters (0.1% of the model) and time (20 minutes). In experiments, RoboMamba demonstrates outstanding reasoning capabilities on general and robotic evaluation benchmarks. Meanwhile, our model showcases impressive pose prediction results in both simulation and real-world experiments, achieving inference speeds 7 times faster than existing robot MLLMs. Our project web page: https://sites.google.com/view/robomamba-web
1 Introduction
The scaling up of data has significantly propelled research on Large Language Models (LLMs) [1, 2, 3], showcasing notable advancements in reasoning and generalization abilities within Natural Language Processing (NLP). To comprehend multimodal information, Multimodal Large Language Models (MLLMs) [4, 5, 6, 7, 8] have been introduced, empowering LLMs with the capability of visual instruction-following and scene understanding. Inspired by the strong capabilities of MLLMs in general settings, recent research aims to incorporate MLLMs into robot manipulation. On one hand, some works [9, 10, 11, 12] enable robots to comprehend natural language and visual scenes, automatically generating task plans. On the other hand, [13, 14, 15] effectively leverage the inherent capabilities of MLLMs, empowering them with the ability to predict manipulation poses.
Robot manipulation involves interacting with objects in dynamic environments, requiring human-like reasoning abilities to comprehend the semantic information of scenes [11, 16], alongside a robust low-level action prediction ability [17, 18]. While existing MLLM-based approaches can handle a range of basic tasks, they still face challenges in two aspects. First, the reasoning capabilities of pre-trained MLLMs [6, 19] in robotic scenarios are found to be insufficient. As shown in Figure 1 (reasoning example), this deficiency presents challenges for fine-tuned robot MLLMs when they encounter complex reasoning tasks. Second, fine-tuning MLLMs and using them to generate robot manipulation actions incurs higher computational costs due to their expensive attention-based LLMs [20, 21]. To balance the reasoning ability and efficiency, several studies [22, 23, 24] have emerged in the field of NLP. Notably, Mamba [25] introduces the innovative selective State Space Model (SSM), promoting context-aware reasoning while maintaining linear complexity. Drawing inspiration from this, we raise a question: “Can we develop an efficient robot MLLM that possesses strong reasoning capabilities while also acquiring robot manipulation skills in a very cost-effective manner?"
To address this, we propose RoboMamba, an end-to-end robotic MLLM that fully leverages the efficiency of Mamba to achieve robust robotic reasoning and action capabilities. As shown in Figure 1, we initially integrate a vision encoder (e.g., CLIP [26]) with Mamba to empower RoboMamba with visual common sense and robot-related reasoning. Specifically, we proceed with alignment pre-training, activating the cross-modal connector [4, 19] to convert visual information into Mamba’s token embeddings. We then unfreeze Mamba for instructions co-training, utilizing its powerful sequence modeling to comprehend high-level robotic and general instruction data. On top of this, to equip RoboMamba with action pose prediction abilities, we explore an efficient fine-tuning strategy with a simple policy head. Notably, we discover that once RoboMamba possesses sufficient reasoning capabilities, it can acquire pose prediction skills with minimal parameter fine-tuning. The fine-tuned policy head constitutes only 0.1% of the model parameters, which is 10 times smaller than existing robot MLLM [15, 14]. In this way, RoboMamba can simultaneously generate robot reasoning using language responses and predict end-effector poses via the policy head.
To systematically evaluate our end-to-end RoboMamba, we conduct extensive experiments in both simulation and real-world scenarios. First, we assess our reasoning abilities on general and robotic evaluation benchmarks. As shown in Figure 1, RoboMamba, with only 3.2B parameters, achieves competitive performance on several MLLM benchmarks and also delivers promising results on RoboVQA (36.3 BLEU-4) [27]. With its strong reasoning abilities, RoboMamba achieves state-of-the-art (SOTA) manipulation performance in the SAPIEN simulation [28], requiring only a 7MB policy head and less than 20 minutes of fine-tuning on a single A100 GPU. Moreover, RoboMamba achieves an inference speed that is 7 times faster than previous SOTA robot MLLM [15]. Additionally, we evaluate RoboMamba in real-world scenarios, where it can generate long-horizon planning and predict the end-effector pose for each atomic task. In summary, our contributions are as follows:
We innovatively integrate a vision encoder with the efficient Mamba language model to construct our end-to-end RoboMamba, which possesses visual common sense and robot-related reasoning abilities.
To equip RoboMamba with action pose prediction abilities, we explore an efficient fine-tuning strategy using a simple policy head. We find that once RoboMamba achieves sufficient reasoning capabilities, it can acquire pose prediction skills with minimal cost.
In our extensive experiments, RoboMamba excels in reasoning on general and robotic evaluation benchmarks, and showcases impressive pose prediction results in both simulation and real-world experiments.
2 Related work
State Space Models (SSMs). SSMs have become effective substitutes for transformers and CNNs due to their linear scalability with sequence length [29, 23]. Recent works [22, 30, 31] use the state space to robustly establish dependencies across long sequences. Especially, Mamba [25] designs the SSM matrices to be functions of the input, creating a learnable selection mechanism that improves adaptability and reasoning capabilities. [32, 33, 34, 35, 36] expand selective SSMs to vision and video tasks. Furthermore, MambaIR [37] focuses on image restoration, and PanMamba [38] addresses pan-sharpening, while DiS [39] integrates SSMs into diffusion models. These findings demonstrate that Mamba exhibits promising performance and efficiency in various visual downstream tasks. With the emergence of SSMs, we make the first attempt to introduce Mamba to address critical challenges in robotics, which demands efficient action capabilities.
Robot Manipulation. Traditional robotic manipulation employs state-based reinforcement learning [40, 41, 42, 43]. In contrast, [44, 11, 45, 46, 47] use state with visual observation as input and then make predictions. Specifically, Where2Act [48] takes visual observations and predicts on actionable pixels and movable regions in objects. Flowbot3d [44] predicts point-wise motion flow on 3D objects. Anygrasp [17] employs point cloud data to learn grasp poses on a large scale datasets. Inspired by the success of MLLMs in general scenarios [49, 50, 19, 51, 52], several studies [13, 16] explore utilizing their common sense reasoning capabilities to address manipulation problems. Palm-E [10] integrates multimodal encodings with LLMs, training them end-to-end for manipulation planning. VoxPoser [11] extracts affordances and constraints from MLLMs to further zero-shot predict trajectories. RoboFlamingo [14] fine-tunes MLLM on vision language manipulation dataset to complete language-conditioned manipulation tasks. ManipLLM [15] introduces specific training scheme for manipulation tasks that equips MLLMs with the ability to predict end-effector poses. ManipVQA [53], enhancing robotic manipulation with physically grounded information processed by MLLM. In this paper, instead of fine-tuning a pre-trained MLLM, we introduce a novel end-to-end MLLM that possesses both robot-related reasoning and pose prediction capabilities. Finally, due to space limitation, we provide additional related work in Appendix F.
3 RoboMamba
In Section 3.1, we introduce the preliminaries of our proposed robot MLLM, including the problem statement and a description of the language model. Subsequently, in Section 3.2 and 3.3, we describe the architecture of RoboMamba and how we empower it with common sense and robot-related reasoning. In Section 3.4, we outline our proposed robot manipulation fine-tuning strategy, which equips our model with pose prediction skills by minimal fine-tuning parameters and time.
3.1 Preliminaries
Problem statement. For robot visual reasoning, our RoboMamba generates a language answer based on the image and the language question , denoted as . The reasoning answer usually contains individual sub-tasks () for one problem . For example, when faced with a planning question like ’How to clean the table?’, the response typically includes steps such as ’Step 1: Pick up the object’ and ’Step 2: Place the object in the box’. For action prediction, we utilize an efficient and simple policy head to predict an action . Following previous works [54, 15], we use 6-DoF to express the end-effector pose of the Franka Emika Panda robot arm. The 6-DoF includes the end-effector position representing a 3D coordinate and direction representing a rotation matrix. If training for a grasping task, we add gripper status to the pose prediction, resulting in a 7-DoF control.
State Space Models (SSMs). In this paper, we select Mamba [25] as our language model. Mamba consists of numerous Mamba blocks, with the most crucial component being the SSM. SSMs [21] are designed based on continuous systems, projecting the 1D input sequence into a 1D output sequence through a hidden state . An SSM consists of three key parameters: the state matrix , the input matrix , and the output matrix . The SSM can be represented as follows:
Recent SSMs (e.g., Mamba [25]) are constructed as discretized continuous systems using a timescale parameter . This parameter transforms the continuous parameters and into their discrete counterparts and . The discretization employs the zero-order hold method, defined as follows:
Mamba introduces the Selective Scan Mechanism (S6) to form its SSM operator in each Mamba block. The SSM parameters are updated to , , and , achieving better content-aware reasoning. The details of the Mamba block are shown in Figure 2.
3.2 RoboMamba architecture
To equip RoboMamba with both visual reasoning and manipulation abilities, we start from pre-trained Large Language Models (LLMs) [25] and visual models to construct an effective MLLM architecture. As shown in Figure 2, we utilize the CLIP visual encoder [26] to extract visual features from input images , where and represent batch size and tokens, respectively. In contrast to [55, 56], we do not adopt the vision encoder ensemble technique, which employs various backbones (i.e., DINOv2 [57], CLIP-ConvNeXt [58], CLIP-ViT) for image feature extraction. The ensemble introduces additional computational costs that severely impact the practicality of robot MLLMs in the real world. Therefore, we demonstrate that a simple and straightforward model design can also achieve strong reasoning abilities when combined with high-quality data and appropriate training strategies. To enable the LLM to understand visual features, we connect the vision encoder to the LLM using a multilayer perceptron (MLP). Through this simple cross-modal connector, RoboMamba can convert visual information into language embedding space . Note that model efficiency is crucial in the field of robotics, as robots need to respond quickly based on human instructions. Therefore, we select Mamba as our language model due to its context-aware reasoning ability and linear computational complexity. Text prompts are encoded into embedding space using the pre-trained tokenizer, then concatenated () with visual tokens and input into Mamba. We leverage Mamba’s powerful sequence modeling to comprehend multimodal information and utilize effective training strategies to develop visual reasoning capabilities (as described in the next section). The output tokens are then detokenized () to produce responses in natural language . The model forward can be represented as follows:
To equip our model with both reasoning and manipulation abilities, we meticulously design a comprehensive training pipeline, which is divided into two stages. We introduce the training recipes of Stage 1 in Section 3.3 and present the robot manipulation fine-tuning in Section 3.4.
3.3 General and robot-related training
After constructing the RoboMamba architecture, the next goal is to train our model to learn general vision and robot-related reasoning abilities. As shown in Figure 2, we divide our Stage 1 training into two steps: alignment pre-training (Stage 1.1) and instruction co-training (Stage 1.2). Specifically, unlike previous MLLM training methods [19, 59, 55], we aim to enable RoboMamba to comprehend both common vision and robotic scenes. Given that the robotics field involves numerous complex and novel tasks, RoboMamba requires enhanced generalization capabilities. Therefore, we adopt a co-training strategy in Stage 1.2, combining high-level robotic data (e.g., task planning) with general instruction data. We find that co-training not only leads to more generalizable robot policies but also enhances general scene reasoning abilities due to the complex reasoning tasks embedded in the robotic data (demonstrated in Appendix C). The training details are shown below:
Stage 1.1: Alignment pre-training. We adopt LLaVA [4] filtered 558k image-text paired dataset for our cross-modal alignment. As shown in Figure 2, we freeze the parameters of the CLIP encoder and Mamba language model, and only update the project layer. In this way, we can align image features with the pre-trained Mamba word embedding.
State 1.2: Instruction co-training. In this stage, we first follow previous MLLM works [4, 55, 56] for general vision instruction data collection. We adopt the 655K LLaVA mixed instruction dataset [4] and 400K LRV-Instruct dataset [60], which aim to learn visual instruction following and mitigate hallucination, respectively. Note that mitigating hallucination plays an important role in robotic scenarios, as the robot MLLM needs to generate task planning based on real scenes instead of imagined ones. For example, existing MLLMs might formulaically answer “open the microwave” with “step 1: find the handle,” but many microwaves do not have handles. Next, we incorporate the 800K RoboVQA dataset [27] to learn high-level robotic skills, such as long-horizon planning, success classification, discriminative and generative affordance, past description, and future prediction. During co-training, as shown in Figure 2, we freeze the parameters of the CLIP encoder and fine-tune the projection layer and Mamba on the combined 1.8 million datasets. All outputs from the Mamba language model are supervised using the cross-entropy loss.
3.4 Robot manipulation fine-tuning
Building upon RoboMamba’s strong reasoning ability, we introduce our robot manipulation fine-tuning strategy in this section, termed Training Stage 2 in Figure 2. Existing MLLM-based manipulation methods [15, 14] require updating the projection layer and the entire LLM during the manipulation fine-tuning stage. While this paradigm can develop action prediction capabilities, it also breaks the inherent abilities of the MLLM and demands significant training resources. To address these challenges, we propose an efficient fine-tuning strategy, as shown in Figure 2. We freeze all the parameters of RoboMamba and introduce a simple policy head to model Mamba’s output tokens. The policy head contains two MLPs that separately learn the end-effector’s position and direction , collectively occupying 0.1% of the entire model parameters. Following [48], the position and direction losses are formulated as follows:
where represents the number of training samples, means the trace of matrix . RoboMamba only predicts the 2D position (, ) of the contact pixel in the image, which is then translated into 3D space using depth information. To evaluate this fine-tuning strategy, we generate a dataset of 10k end-effector pose predictions using the SAPIEN simulation [28]. After manipulation fine-tuning, we find that once RoboMamba possesses sufficient reasoning capabilities, it can acquire pose prediction skills with extremely efficient fine-tuning. Due to the minimal fine-tuning parameters (7MB) and efficient model design, we need only 20 minutes to achieve novel manipulation skills. This finding highlights the importance of reasoning abilities for learning manipulation skills and presents a new perspective: we can efficiently equip an MLLM with manipulation abilities without compromising its inherent reasoning capabilities. Finally, RoboMamba can use language responses for common sense and robot-related reasoning, and the policy head for action pose prediction.
4 Experiment
In Section 4.1, we introduce our experiment settings, including dataset, implementation, and evaluation benchmark details. Subsequently, we conduct extensive experiments to demonstrate RoboMamba’s reasoning and manipulation abilities in Sections 4.2 and 4.3, respectively. To thoroughly validate the effectiveness of each method design, we perform an ablation study in Section 4.4. Finally, the qualitative results of real-world experiments are presented in Section 4.5.
4.1 Experiment settings
Datasets (Stage 1) In the alignment pre-training stage, we utilize the LLaVA-LCS 558K dataset [59], which is a curated subset of the LAION-CC-SBU dataset, supplemented with captions. During the instruction co-training stage, we created a combined dataset totaling 1.8 million samples, including LLaVA-v1.5 655K mix [59], LRV-INSTRUCT 400K [60], and RoboVQA 800K dataset [27]. The detailed descriptions of the datasets is shown in Appendix B.
Datasets (Stage 2) For the dataset used in the robot manipulation fine-tuning stage, we utilize the SAPIEN [28] to set up an interactive simulation environment with articulated objects from PartNet [45]. To generate data, we use the Franka Panda Robot to randomly interact with objects. When successful manipulation occurs, we record the success 6-DOF poses of the end-effector, which serves as ground truth labels for training. In the training set, we collect 10K images across 20 categories. For testing, we use a set of 1.1K images that include both seen categories from the training set and unseen categories. The details of the categories are provided in Appendix B.
Implementation details Before training, RoboMamba loads a pre-trained CLIP/SigLIP ViT-Large [26, 61] as the visual encoder, and the 2.8/1.4B Mamba [1] model as the language model. During the alignment pre-training and instruction co-training, we conduct training for 1 epoch and 2 epochs, respectively. We utilize the AdamW optimizer with and a learning rate (LR) of 2e-5. The precision of floating-point calculations is set to 16-bit. For manipulation fine-tuning, we train the model for 5 epochs, setting the LR to 1e-5 and applying a weight decay of 0.1. The floating-point precision is set to 32-bit. All experiments are conducted on NVIDIA A100 GPUs.
Reasoning evaluation benchmarks To evaluate reasoning capabilities, we employ several popular benchmarks, including VQAv2 [62], OKVQA [63], RoboVQA [27], GQA [64], OCRVQA [65], VizWiz [66], POPE [67], MME [68], MMBench [69], and MM-Vet [70]. As detailed in Appendix E, we describe the key aspects each benchmark focuses on when assessing models in the field of robotics. Notably, we also directly evaluate RoboMamba’s robot-related reasoning abilities on the 18k validation dataset of RoboVQA, covering robotic tasks such as long-horizon planning, success classification, discriminative and generative affordance, past description, and future prediction.
Manipulation evaluation benchmarks To evaluate our model’s manipulation capabilities, we follow previous works [44, 54, 15] and test pulling accuracy exclusively in the simulator [28]. We use the predicted contact point and rotation to interact with objects. To measure the model’s performance, we use the classical manipulation success rate, defined as the ratio of successfully manipulated samples to the total test samples. A manipulation action is considered successful if the difference in the object’s joint state before and after interaction exceeds a threshold of 0.1 meters. In real-world experiments, we use the Franka Panda robot to manipulate several articulated objects.
Method LLM Size Res. OKVQA VQAV2 GQA VizWiz OCR-VQA POPE MME MMB MM-Vet BLIP-2 [49] 7B 224 45.9 - 41.0 19.6 40.6 85.3 1293.8 - 22.4 InstructBLIP [71] 7B 224 - - 49.5 33.4 44.8 - - 36 26.2 LLaMA-AdapterV2 [51] 7B 336 49.6 70.7 45.1 39.8 - - 1328.4 - - MiniGPT-v2 [72] 7B 448 57.8 - 60.1 53.6 - - - - - Qwen-VL [73] 7B 448 58.6 79.5 59.3 35.2 75.7 - - 38.2 - LLaVA1.5 [59] 7B 336 - 78.5 62.0 50.0 - 85.9 1510.7 64.3 30.5 SPHINX [55] 7B 224 62.1 78.1 62.6 39.9 66.0 80.7 1476.1 66.9 36.0 LLaVA-Phi [74] 2.7B 336 - 71.4 - 35.9 - 85.0 1335.1 59.8 28.9 MobileVLM [75] 2.7B 336 - - 59.0 - - 84.9 1288.9 59.6 - TinyLLaVA [76] 2.7B 336 - 77.7 61.0 - - 86.3 1437.3 68.3 31.7 RoboMamba(Ours) 2.7B 224 63.1 80.3 62.4 55.0 62.5 85.3 1314.8 64.2 28.6 RoboMamba(Ours) 2.7B 384 62.4 79.1 64.4 55.0 66.7 86.9 1354.2 65.7 29.7
4.2 Reasoning quantitative result
General reasoning. As shown in Table 1, we compare RoboMamba with previous state-of-the-art (SOTA) MLLMs on general VQA and recent MLLM benchmarks. First, we find that RoboMamba achieves promising results across all VQA benchmarks, using only a 2.7B language model. The results demonstrate that our simple architecture design is effective. The alignment pre-training and proposed instruction co-training significantly enhance the MLLM’s reasoning capabilities. For example, due to the large amount of robot data introduced during the co-training stage, our model’s spatial identification performance on the GQA benchmark is improved. Meanwhile, we also test our RoboMamba on recently proposed MLLM benchmarks. Compared to previous MLLMs, we observe that our model achieves competitive results across all benchmarks. Notably, our model achieves satisfactory results on the POPE benchmark due to the inclusion of the LRV-Instruct dataset during the co-training stage, which helps reduce failed robot actions caused by hallucinations. Although some performances of RoboMamba are still below those of LLaVA1.5 and SPHINX, we prioritize using a smaller and faster Mamba to balance the efficiency of the robotic model. In the future, we plan to develop RoboMamba-7B for scenarios where resources are not limited.
Robot-related reasoning. To comprehensively compare RoboMamba’s robot-related reasoning abilities, we benchmark it against LLaMA-AdapterV2 [51] on the RoboVQA [27] validation set. We chose LLaMA-AdapterV2 as a baseline because it serves as the base model for the current SOTA Robot MLLM, ManipLLM [15]. For a fair comparison, we loaded the baseline pre-trained parameters and fine-tuned it on the RoboVQA training set for two epochs, using its official instruction tuning method. As shown in Figure 3 a), RoboMamba achieves superior performance across BLEU-1 to BLEU-4. The results indicate that our model possesses advanced robot-related reasoning capabilities and confirms the effectiveness of our training strategy. In addition to higher accuracy, our model achieves inference speeds 7 times faster than LLaMA-AdapterV2 and ManipLLM, which can be attributed to the content-aware reasoning ability and efficiency of the Mamba language model [25]. Finally, we visualize the qualitative results in Figure 4.
| Seen Categories | ||||||||||||||||
| Method | ||||||||||||||||
| UMPNet [54] | 0.28 | 0.41 | 0.25 | 0.20 | 0.49 | 0.20 | 0.35 | 0.57 | 0.51 | 0.25 | 0.66 | 0.17 | 0.17 | 0.26 | 0.27 | 0.40 |
| FlowBot3D [44] | 0.50 | 0.53 | 0.26 | 0.36 | 0.34 | 0.36 | 0.54 | 0.26 | 0.12 | 0.34 | 0.41 | 0.23 | 0.36 | 0.30 | 0.17 | 0.37 |
| RoboFlamingo [14] | 0.48 | 0.51 | 0.50 | 0.35 | 0.11 | 0.47 | 0.54 | 0.35 | 0.19 | 0.46 | 0.18 | 0.64 | 0.26 | 0.42 | 0.15 | 0.87 |
| ManipLLM [15] | 0.68 | 0.62 | 0.45 | 0.74 | 0.42 | 0.25 | 0.61 | 0.66 | 0.56 | 0.52 | 0.50 | 0.42 | 0.64 | 0.76 | 0.63 | 0.60 |
| RoboMamba (Ours) | 0.81 | 0.73 | 0.33 | 0.85 | 0.86 | 0.60 | 0.81 | 0.42 | 0.56 | 0.54 | 0.68 | 0.81 | 0.26 | 0.86 | 0.39 | 0.91 |
| Seen Categories | Unseen Categories | |||||||||||||||
| Method | AVG | AVG | ||||||||||||||
| UMPNet [54] | 0.27 | 0.37 | 0.19 | 0.60 | 0.34 | 0.32 | 0.36 | 0.18 | 0.37 | 0.21 | 0.12 | 0.04 | 0.53 | 0.28 | 0.13 | 0.26 |
| FlowBot3D [44] | 0.21 | 0.57 | 0.29 | 0.45 | 0.35 | 0.36 | 0.36 | 0.18 | 0.30 | 0.21 | 0.50 | 0.13 | 0.53 | 0.28 | 0.09 | 0.30 |
| RoboFlamingo [14] | 0.20 | 0.42 | 0.58 | 0.60 | 0.41 | 0.36 | 0.62 | 0.64 | 0.33 | 0.14 | 0.34 | 0.44 | 0.66 | 0.41 | 0.31 | 0.43 |
| ManipLLM [15] | 0.41 | 0.78 | 0.41 | 0.59 | 0.56 | 0.21 | 0.25 | 0.79 | 0.76 | 0.52 | 0.76 | 0.43 | 0.85 | 0.26 | 0.52 | 0.51 |
| RoboMamba(Ours) | 0.40 | 0.55 | 0.37 | 0.80 | 0.63 | 0.19 | 0.23 | 0.67 | 0.66 | 0.57 | 0.45 | 0.65 | 0.68 | 0.30 | 0.93 | 0.53 |
4.3 Manipulation quantitative result
Baselines. To evaluate RoboMamba’s manipulation abilities, we compare our model with four baselines: UMPNet [54], Flowbot3D [44], RoboFlamingo [14], and ManipLLM [14]. Before comparison, we reproduce all baselines and train them on our collected dataset. For UMPNet, we execute manipulation on the predicted contact point, with the orientation perpendicular to the object’s surface. Flowbot3D predicts motion direction on the point cloud, selecting the largest flow magnitude as the interaction point and using the direction of the flow to represent the end-effector’s orientation. RoboFlamingo and ManipLLM separately load the pre-trained parameters of OpenFlamingo [50] and LLaMA-AdapterV2 [51], and follow their respective fine-tuning and model updating strategies.
Results. As shown in Table 2, our RoboMamba achieves a 7.0% improvement on seen categories and a 2.0% improvement on unseen categories compared to the previous SOTA ManipLLM. Moreover, our method showcases SOTA performance across 14 of 20 seen categories, highlighting its effectiveness and stability in predicting action poses. For unseen categories, the recent three MLLM-based methods—RoboFlamingo, ManipLLM, and our method—all achieved promising performance. The results demonstrate that leveraging the strong generalization abilities of MLLMs can effectively improve the policy’s generalization ability while enhancing accuracy on unseen objects. Regarding efficiency, RoboFlamingo updates 35.5% (1.8B) of the model parameters, ManipLLM updates an adapter (41.3M) comprising 0.5% of the model parameters, whereas our fine-tuned simple policy head (3.7M) only constitutes 0.1% of the model parameters. RoboMamba effectively updates 10 times fewer parameters than previous MLLM-based methods while achieving seven times faster inference speeds. The results reveal that our RoboMamba not only possesses strong reasoning abilities but also can acquires manipulation capabilities in a cost-effective manner.
4.4 Ablation study
The impact of reasoning ability. We explore whether utilizing MLLMs with different reasoning abilities affects manipulation skill learning. For a fair comparison, we use the same manipulation fine-tuning strategy, injecting and fine-tuning a simple MLP policy head after the MLLM (while freezing other parameters). We compare our RoboMamba-2.7B (Ours-2.7B) with OpenFlamingo, LLaMA-AdapterV2, and our RoboMamba-1.4B. As shown in Figure 3 b), Ours-2.7B achieves promising results compared with other methods, which is proportional to its reasoning ability. Meanwhile, Ours-2.7B (w/o C) indicates that we did not use the instruction co-training method, omitting the robot-related RoboVQA dataset during fine-tuning. We find that this also impacts the accuracy of manipulation, especially reducing the model’s generalization ability when facing unseen objects. The results confirm our finding: fine-tuning an MLLM to learn robot skills does not require extensive resources; it only requires that the MLLM possesses strong robot-related reasoning abilities. Additionally, we present more ablation studies in Appendix C, including explorations of training strategies and policy head design.
4.5 Real-world experiments
As shown in Figure 4, we visualize RoboMamba’s reasoning results across various robotic downstream tasks. For task planning, compared to LLaMA-AdapterV2, RoboMamba demonstrates more accurate and long-horizon planning abilities, thanks to its strong reasoning capabilities. For a fair comparison, we also fine-tuned the baseline LLaMA-AdapterV2 on the RoboVQA dataset. Additionally, RoboMamba accurately performs fundamental robotic tasks such as affordance generation and discrimination, proving that it can understand robotic scenes. Notably, our model also possesses past and future prediction capabilities, further highlighting its robust reasoning capabilities. For pose prediction, we use a Franka Emika robotic arm to interact with various household objects. We project RoboMamba’s predicted 3D pose back onto a 2D image, using a red dot to indicate the contact point and the end-effector to show the direction, as shown in the bottom right corner of the figure. More real-world demonstrations are provided in Appendix 5 and the supplementary video file.
5 Conclusion and future plan
In this paper, we introduce an end-to-end robotic MLLM named RoboMamba, which possesses both reasoning and manipulation capabilities. Based on our RoboMamba, we can impart new manipulation skills to the model by fine-tuning a simple policy head (0.1% of the model) in approximately 20 minutes. This finding reveals how to efficiently equip an MLLM with manipulation abilities without compromising its inherent reasoning capabilities. Finally, RoboMamba excels in reasoning on both general and robot-related evaluation benchmarks and showcases impressive pose prediction results. As for our future plan, we focus on two directions. 1) During the process for robot manipulation fine-tuning, we aim to introduce continual learning techniques [77, 78, 79] to continuously enhance the model’s manipulation abilities in the real world. 2) Constructing a 3D Robot MLLM [80, 81, 82], as the 3D point cloud contains more geometric information that helps predict 3D manipulation poses.
Appendix A Appendix
Due to space limitations, we provide additional details of the proposed method in this supplementary material. In Appendix B, we offer a more detailed description of our training dataset, including alignment pre-training, instruction co-training, and robot manipulation fine-tuning. Additional ablation study are presented in Appendix C, which explore the impact of different training strategies on reasoning ability and the effect of different head designs on manipulation fine-tuning. In Appendix 5, we show additional qualitative results across multiple robot-related downstream tasks. Finally, we provide the metric selection rationale and the usage of prompts during testing in Appendix E.
Appendix B Dataset description
| Safe | Door | Display | Refrigerator | Laptop | Lighter | Microwave | Mouse | Box | Trashcan |
| Kitchen pot | Suitcase | Pliers | Storage | Remote | Bottle | Folding chair | Toaster | Lamp | Dispenser |
| Toilet | Scissors | Table | Stapler | Kettle | USB | Switch | Washing | Faucet | Phone |
Stage 1.1: Alignment pre-training dataset
1) LLaVA-LCS 558K: This LLaVA Visual Instruct Pretrain LCS-558K dataset is a curated subset of the LAION/CC/SBU dataset, specifically filtered to achieve a more balanced distribution of concept coverage. Additionally, it includes captions paired with BLIP synthetic captions for reference purposes.
Stage 1.2: Instruction co-training dataset.
1) LLaVA-v1.5 655K: This dataset is a mixture of ten distinct datasets, including LLaVA [4], ShareGPT [83], VQAv2 [62], GQA [64], OKVQA [63], OCRVQA [63], A-OKVQA [84], TextCaps [85], RefCOCO [86, 87], and Visual Genome (VG) [88]. This mix dataset is also one of the most renowned datasets used for instruction tuning in several works [4, 59].
2) LRV-INSTRUCT 400K: Although recent work [55] suggests that increasing image resolution can effectively mitigate hallucination, processing high-resolution images affects the efficiency of robotic policy. Therefore, we introduced this dataset during the co-training process to mitigate hallucination. This dataset contains visual instructions generated by GPT-4, encompassing 16 different vision-and-language tasks with open-ended instructions and answers.
3) RoboVQA 800K: In co-training, we use this dataset to enhance our model’s robot-related reasoning abilities. RoboVQA [27] comprises realistic data collected by performing various user requests and using multiple embodiments, such as robots, humans, and humans with grasping tools. This dataset includes 5,246 long-horizon episodes and 92,948 medium-horizon episodes of robotic tasks, each paired with image and text prompt inputs.
Stage 2: Robot manipulation fine-tuning dataset.
Representation for Each Category Icon In Table 3, we provide an overview of the meaning of each category icon presented in Table 2 of the main paper. These categories and their corresponding objects are sourced from PartNet-Mobility [89].
Simulator Data Collection In the simulator, we use a Franka Panda Robot with a suction gripper as the robotic actuator. During data collection, we randomly select a contact point on the movable part of the object and orient the end-effector’s z-axis opposite to the object’s normal vector, with a random y-axis direction to interact with the object. Successful operations are categorized as successful samples and integrated into the training dataset. For the training set, we collect 10K images across 20 categories, including Safe, Door, Display, Refrigerator, Laptop, Lighter, Microwave, Mouse, Box, Trash Can, Kitchen Pot, Suitcase, Pliers, Storage Furniture, Remote, Bottle, Folding Chair, Toaster, Lamp, and Dispenser. For testing, we use a set of 1.1K images that include both seen categories from training and unseen categories, such as Toilet, Scissors, Table, Stapler, Kettle, USB, Switch, Washing Machine, Faucet, and Phone.
Appendix C Additional ablation study
The impact of training strategies on reasoning abilities As shown in Table 4, we explore the impact of different training strategies on common sense reasoning ability. Specifically, we conduct these experiments using input images. In this table, AP refers to alignment pre-training, and IC refers to instruction co-training with different dataset combinations. First, we observe that Ex2 outperforms Ex1 across all three metrics, validating the importance of alignment pre-training, consistent with conclusions drawn by mainstream methods [19, 4, 49]. Next, we find that incorporating the LRV-INSTRUCT 400K dataset (Ex3) indeed improves POPE accuracy, helping our model mitigate the negative effects of hallucinations. Finally, introducing robot-related datasets in co-training not only empowers our model with robot-related reasoning abilities but also enhances common sense reasoning performance, particularly on benchmarks like GQA that are related to spatial reasoning.
| AP | IC(LLaVA-655K) | IC(LRV-400k) | IC(Robo-800k) | OKVQA | GQA | POPE | |
| Ex1 | - | ✓ | - | - | 61.5 | 62.2 | 85.5 |
| Ex2 | ✓ | ✓ | - | - | 62.3 | 62.7 | 85.9 |
| Ex3 | ✓ | ✓ | ✓ | - | 62.0 | 62.6 | 86.6 |
| Ex4 | ✓ | ✓ | ✓ | ✓ | 62.4 | 63.8 | 86.9 |
The impact of policy head designs on manipulation accuracy As shown in Table 5, we explore the impact of different policy head designs on manipulation skill learning. In this table, MLP1 means using only one MLP heads to predict the position and direction of the end-effector pose. MLP2 means using one shared head to predict direction and another head to predict position separately. (SSM block+MLP)2 is similar to MLP2 but adds a State Space Model (SSM) block before the MLP to increase the parameter count of the policy head. The experimental results show that the manipulation accuracy across the three configurations is quite similar, indicating that the parameter count of the fine-tuning policy head has small impact on the results. Combined with Figure 3 b), this further supports our finding that once RoboMamba achieves sufficient robotic reasoning capabilities, it can acquire pose prediction skills at a low cost, regardless of the policy head design.
| result | MLP2 | MLP1 | (SSM block+MLP)2 |
| Acc (Seen) | 63.7% | 62.1% | 63.2% |
| Parameters | 3.7M | 1.8M | 45.2M |
| Percentage | 0.11% | 0.05% | 1.3% |
Appendix D Additional real-world experiments
We conduct real-world experiments involving interactions with various household objects using a Franka Emika robotic arm. We modify the finger gripper by attaching double-sided tape to convert it into a suction gripper, providing the gripper head with adhesive properties. The video demonstrations are included in the supplementary video file. As shown in Figure 5, we visualize our model’s reasoning results on a series of robotic downstream tasks, including long-horizon planning, discriminative affordance, generative affordance, past description, and future prediction. Additionally, we project our RoboMamba’s predicted 3D pose back onto a 2D image based on camera parameter.
Appendix E Reasoning evaluation bencharks
To comprehensively evaluate our model’s capabilities in common sense reasoning, we select several general MLLM evaluation benchmarks, prioritizing those related to robotics. Below is a description of each benchmark
VQAv2 and OKVQA: These benchmarks are utilized to assess the model’s proficiency in basic vision question answering, which is a foundational skill in embodied AI. This ability ensures that the model can understand and respond to visual content effectively.
POPE and VizWiz: These benchmarks are chosen to evaluate the model’s capability to answer questions without falling prey to visual illusions or ambiguities. This aspect is crucial for avoiding significant errors in robotic applications.
GQA and OCRVQA: These benchmarks are employed to test the model’s ability to identify and comprehend the types and positions of important objects within an image. Such spatial identification skills are vital for tasks related to robotic manipulation and interaction with the environment.
RobotVQA: This benchmark is used to assess the model’s ability to plan and understand actions based on both textual and visual inputs. This skill is indispensable in the realm of robotics, where understanding and executing complex actions is necessary.
MM-Vet, MME and MMB: These benchmarks are utilized to evaluate multimodal large language models’s ability to integrate on complex multi-modal tasks including Recognition, Spatial awareness, OCR, and Math. All of them contain a wealth of evaluation indicators, such as perception and cognition, which can fully demonstrate the performance of the model under different tasks, and this performance is the best embodiment of the comprehensive application performance of multimodal large language models(MLLM).
Appendix F Additional related work
Multimodal Large Language Models. Large language models (LLMs) have exhibited remarkable reasoning capabilities across various downstream tasks [19, 90, 2]. When addressing complex multimodal reasoning challenges, multimodal large language models (MLLMs) have shown exceptional visual understanding, i.e., BLIP-2 [49], OpenFlamingo [50], LLaMA-Adapter [19, 51], and LLaVA [52]. Additionally, the introduction of 3D MLLMs [12, 91, 81] seeks to expand the reasoning and conversational capabilities of LLMs to include the 3D modality. However, deploying LMMs is expensive due to their significant computational overhead, primarily caused by their billions of parameters. To mitigate these challenges, recent small-scale models [74, 92] demonstrate impressive performance while maintaining manageable computational costs. LLaVA-Phi [74] empowers the recently developed smaller LLM, Phi-2, for visual instruction tuning. TinyLLaVA [92] and MobileVLM V2 [93] demonstrate that high-quality training data and schemes can effectively compensate for the reasoning abilities of smaller LMMs. Furthermore, Cobra [56] innovatively utilizes an SSM-based language model to reduce complexity. Different from previous works, our goal is to develop an efficient embodied MLLM using the Mamba language model. This model not only possesses common sense understanding but also has the capability to complete manipulation tasks effectively.